A certain amount is invested at 9% and 12,000 more than that amount is invested at 10%. How much is invested at each rate if the total income is 10,130?
Answer by nerdybill(7384)

A certain amount is invested at 9% and 12,000 more than that amount is invested at 10%. How much is invested at each rate if the total income is 10,130?
.
Let x = amount ($) invested at 9%
then
x+12000 = amount ($) invested at 10%
.
.09x + .10(x+12000) = 10130
.09x + .10x+1200 = 10130
.19x+1200 = 10130
.19x = 8930
x = 8930/.19
x = $47000 (invested at 9%)
.
Amount invested at 10%:
x+12000 = 47000+12000 = $59000